OpenStudy (anonymous):

Solve the following system of equations: a)y=-2x-1 b)2x-7y=3 x+2y=4 -5x+3y=7

OpenStudy (anonymous):

A- y=-2x-1 } *2 \[\rightarrow 2y =-4x -2\] -eq 1- 2y=-x+4 -eq2- by subtracting eq1 from eq2 : \[(2y=-4x-2)-(2y=-x+4) \rightarrow 0=-3x-6 \] so: x=-2 then from eq1 \[\rightarrow 2y=-4(-2) -2 = 8-2 =6\] then y=6/2 =3

OpenStudy (anonymous):

is it clear ??